OpenStudy (anonymous):

The areas of two similar triangles are 16cm^2 and 25 cm^2. One of the sides of the second triangle is 8cm. What is the length of the corresponding side of the other triangle?

OpenStudy (anonymous):

@dumbcow please help me!!!

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (dumbcow):

ratio of areas = (ratio of sides)^2 \[\frac{16}{25} = (\frac{x}{8})^{2}\] solve for x \[x = \sqrt{\frac{(64)(16)}{25}} = \frac{(8)(4)}{5} = \frac{32}{5}\]